Andrew Sward

University of Illinois at Chicago, Department of Mathematics, 851 S. Morgan (M/C 249) Chicago, Illinois 60607-7045 (508) 873-4834, asward2@uic.edu http://www.math.uic.edu/∼apsward/

DATE OF BIRTH: December 28, 1984

EDUCATION 2009 MS in Mathematics (expected May 2009) from the University of Illinois at Chicago. 2007 BS in Computer Science and Mathematics at Northeastern University.

TEACHING EXPERIENCE • Teaching Assistant (August 2007)

– Mathematical Reasoning (discussion leader).

– Participated in the Mathematics Tutoring Lab

• Supplemental Instructor (2006-2007) – Differential Equations and Linear Algebra (discussion leader).

– Held office hours for students.

COMPUTER SCIENCE EXPERIENCE – Wrote Ray-Tracing program for spheres in java. (April 2007)

– Co-wrote a logic-based SuDoKu solver in java. (June 2006)

– Junior Software Developer Co-op postion with Empirix (June 2005 - Dec 2005).

– Junior Software Developer Co-op postion with Empirix (June 2004 - Dec 2004).

– Mathematica, Maple, Matlab C, C++, java, scheme.